OpenStudy (anonymous):

Solve the system of linear equations below. 2x + 3y = 2 x + 6y = 4

OpenStudy (anonymous):

2x+3y=2 x+6y=4 To make it more clear.

OpenStudy (cydney_morgan):

We will mark equation 1 as E1 and equation 2 as E2 First we will switch E1 with E2 2x + 3y = 2 x + 6y = 4 ---> x + 6y = 4 2x + 3y = 2 Now will do following E2=E2 - 2*E1 ----> x + 6y =4 ----> x + 6y = 4 (2-2)x + (3-2*6)y =2 -2*4 0*x + -9y = -6 Now we will work only in the second equation -9y =-6 y=6/9 =2/3 We will insert it in 2x + 3y = 2 2x +3 *(2/3) =2 2x +2 =2 /-2 2x=0 x=0 The solution will be (0,2/3) ( Curtsy of Google Because I'm not that smart )

OpenStudy (zzr0ck3r):

2x+3y=2 x+6y=4 as @iambatman said it makes it easy if we write it like this, note if we multiply the second equation by -2 on both sides (which we are allowed to do), we get 2x+3y=2 -2x-12y=-8 now notice the 2x on top will cancel with the -2x on bottom if we add the equations 2x+(-2x)+3y+(-12y)=-6 which becomes -9y=-6 so y= 6/9 now plug this into either equation to find x.